OpenStudy (anonymous):

Two families leave the same spot in a park at the same time. One family drives 6 miles north and the other drives directly west. When they stop for a break, the distance between them is 10 miles. How far did the family driving west travel?

OpenStudy (anonymous):

This question belongs in the mathematics section, just FYI. If you draw a diagram of this problem, it looks like this: From there, you should use the Pythagorean Theorum to find the length of the missing side. Do you know how to do that?

OpenStudy (anonymous):

I know it belongs in the math section, sorry was not really worried about that part. And no I don't.

OpenStudy (anonymous):

Hmm. The pythagorean theorem says that a^2 + b^2 = c^2. C represents the hypotenuse, while a and b are the other 2 sides. So in this case you know a and c, you just need to find b. first, plug in the values of a and c. What does that equation look like?

OpenStudy (anonymous):

(hint: a is 6 and c is 10)

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Not quite. That would be totally true if it were a+b=c. Unfortunately instead it's a squared plus b squared = c squared. So it's 6^2 + b^2 = 10^2 So 36 + b^2 = 100 Now try solving that for b.

OpenStudy (anonymous):

@Tennisstar16

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What did you get when you solved for b?